# Towards a more realistic evaluation of machine learning models for bearing fault diagnosis

### v2

@@ 1 Introduction
Establishing that conventional (non-bearing-wise) splitting strategies indeed lead to unrealistic results—and thus should be avoided in future work—requires one to show that using a bearing-wise split causes a significant performance gap in an otherwise identical setup. This path was taken by [6, 7, 8, 9], which have observed accuracy dropping from near 100% to around 40%-60% depending on the specific setup. However, [7, 8] proposed splitting data by fault size, which does not necessarily correspond to a bearing-wise split (see Section 3.4), while [7, 9] could not entirely eliminate leakage with respect to the healthy class in the CWRU dataset, as this dataset contains a single healthy bearing configuration.
@@
An alternative approach to completely eliminating data leakage is conducting inter-testbench experiments, typically framed within a cross-domain or domain generalization context. In this formulation, the testbenches used for training and testing are completely isolated; however, this introduces the significant challenge of domain generalization. We suspect that the bearing fault diagnosis literature has largely treated intra-domain classification as a solved problem, prompting a shift toward cross-domain research. Research by [10] indicates that many studies define a domain as a combination of operating conditions (e.g., load, rotation speed, torque) rather than distinct physical bearings. This perspective has led to various data splitting methodologies within single datasets, where data is partitioned according to these conditions. A limited number of studies have addressed the inter-testbench scenario by isolating datasets, such as [11], [12], and [13]. Although these approaches effectively eliminate data leakage, domain generalization remains a complex challenge, as target datasets often exhibit disparate feature distributions, necessitating specialized techniques to extract domain-invariant features. The present work, however, focuses on a simpler yet fundamentally critical problem: training and testing within a single test bench.
@@
This paper aims to further advance the reliable development and deployment of ML models for bearing fault diagnosis. We propose a novel methodology that rigorously addresses data leakage and class imbalance in a intra-dataset scenario, particularly for datasets with limited healthy bearing data. Our approach formulates the problem as a binary multi-label classification for each sensor location (e.g., drive end and fan end), enabling the detection of the presence or absence of each fault type (e.g., inner, outer, ball). This formulation specifically addresses the disadvantages of multiclass accuracy, which serves as a poor proxy for real-world performance. Because accuracy treats all misclassifications as equal, it fails to distinguish between “false alarm” (False Positive) and a “missed detection” (False Negative), the latter being significantly more costly in industrial maintenance. Furthermore, in the presence of class imbalance, a model can achieve high accuracy by simply predicting the majority class, effectively “hiding” its inability to identify rare but critical fault states. Another issue caused by the multiclass formulation is the inability of detecting co-occurring faults. While it is possible to create classes for combined faults, it is often unpractical and most public datasets contain few examples of those cases. By treating faults as independent binary problems, we can accommodate co-occurring defects and utilize faulty signals from one label as true negatives for others—a significant advantage when healthy data is scarce, such as in the CWRU dataset. In this formulation, we can also utilize prevalence-independent metrics such as the Area Under the Receiver Operating Characteristic curve (AUROC). This allows for a more realistic representation of real-world conditions where decision thresholds must be precisely tuned to prioritize sensitivity or specificity based on specific safety and operational requirements.
@@
With these contributions, this work aims to foster the development of more robust and trustworthy ML models for bearing fault diagnosis, ensuring their performance more closely reflects their capabilities in real-world industrial settings. The source code for this paper can be found at https://github.com/gama-ufsc/bearing-data-leakage.
@@ 3.2 A Toy Example
To demonstrate the performance inflation caused by a non-bearing-wise data split, we constructed a synthetic binary fault detection experiment, in which the task is to classify each sample as either faulty or healthy. We simulate a feature space for a dataset of $B=48$ distinct bearings, which could be obtained, for example, as the output of a deep learning model’s feature extractor. This simulated space contains two feature types: 3 fault-predictive features and 48 bearing-identity features, totalizing 51 features. Each identity feature is unique to a specific bearing, representing a spurious correlation unrelated to the fault condition. All features were generated by adding zero-mean Gaussian noise with unit variance to a constant base value: $a_{f}=1.5$ for fault-predictive features and $a_{b}=8$ for bearing-identity features. All experiments were designed to maintain a fixed number of samples per bearing, using a base value of 40 samples each. The dataset comprised 24 healthy bearings and 24 faulty bearings, ensuring a consistently balanced class distribution.
@@ 3.3 Prevalence of data leakage in bearing fault diagnosis
As mentioned in the introduction, data leakage is widespread in the literature on bearing fault diagnosis, with previous works [4, 6] reporting data splitting issues in over 90% of published papers. To complement and update these findings, we conducted an investigation of papers published in 2025 using a similar methodology. First, we searched the Mechanical Systems and Signal Processing journal database using the queries “bearing fault diagnosis” and “machine learning”, which yielded 195 published papers from January 2025 up to December 2025. We then randomly selected 10 papers that focus on the problem of training and testing within a single dataset under a supervised machine learning context. Second, we extended our search to include papers from other journals within the same scope, compiling an additional 8 papers.
Table 1 lists these 18 papers, detailing whether they specified the train-test partitioning and, if so, how it was performed. We found that 8 of the 18 papers used a random splitting strategy, while 9 detailed a condition-wise split. The remaining paper did not mention any partitioning methodology. The most common condition considered for partitioning was the load, although other papers considered rotation speed or level of noise. Our conclusion is that despite an increase in the number of articles that detail their train-test partitioning methodology, data leakage remains a prevalent issue.
@@ 3.4 Related works
Hendriks et al. [7] proposed a fault-size splitting strategy for the CWRU dataset, considering it a better approach to obtain domain shift in bearing fault diagnosis. Since in the CWRU dataset each combination of fault size and type (i.e. inner, outer, ball) correspond to a unique bearing, splitting by fault size indirectly produces as a bearing-wise partition. However, signals from the single healthy configuration (consisting of a healthy bearing in the fan end and another healthy bearing in the drive end) were included in both training and test sets, resulting in data leakage.
@@
A comprehensive investigation into dataset biases and evaluation protocols was conducted by [5], who proposed evaluation methodologies for widely used datasets such as CWRU, PU, MFPT, IMS, and UOC. Their approach focused on mitigating segmentation-level leakage, advocating for condition-wise splits across all datasets under a multiclass framework, with F1-macro reported as an auxiliary metric. However, the results remained overoptimistic due to residual bearing-level leakage, which was not addressed in their papers.
Another notable contribution is the study by [6], which proposed a bearing-wise split under a multiclass classification setting, referred to as a “part-to-part” approach. Using the KAt and CMTH datasets, the authors demonstrated a high correlation between signals originating from the same bearing—violating the i.i.d. assumption—and showed the substantial drop in accuracy when shifting from a condition-wise to a bearing-wise split, revealing the effects of a bearing-level leakage. Although this paper propose a similar analysis to ours, their experiments contain confounding factors that make it difficult to understand the origin of the decrease in model performance. In their work, data splitting strategies such as “run-to-run”, “day-to-day” and “part-to-part” were proposed. Each strategy correspond to completely different training set compositions, making it difficult to identify which one is more diverse. One could argue that the decrease in performance when changing to the “part-to-part” split is associated with the model being less diverse, rather than uniquely due to data leakage. Our paper focuses on solving this issue by proposing controlled data leakage experiments, where the training set remains fixed while varying the test sets.
@@ 4.2.2 Paderborn University (PU) Dataset
Crucially, prior work by [41] demonstrated that a significant domain shift exists between artificially damaged and naturally degraded bearings, leading to poor generalization when training on the former and testing on the latter. To circumvent this issue and ensure the practical relevance of our findings, our investigation exclusively utilizes the subset of bearings with naturally occurring faults from accelerated lifetime tests, along with the healthy reference bearings. Our analysis incorporates components K006 and KI17, thereby expanding upon the set used in the original benchmark study, which are shown in Table 3. Although bearings with combined faults are in principle compatible with our multi-label methodology, their limited representation in the dataset hinders robust learning and evaluation. In fact, only three such bearings are available (IDs KB23, KB24, and KB27). Therefore, we exclude these samples from both training and testing to prevent biased or unreliable conclusions.
@@ 4.2.3 CWRU bearing fault dataset
The Case Western Reserve University (CWRU) bearing fault dataset [44] is a collection of experiments that involved a single pair of healthy bearings and several artificially created faulty bearings. The faults were created through electro-discharge machining, introducing point faults with diameters of 7, 14, 21, and 28 mils in the inner race, outer race, and rolling element separately. For the outer race faults, the experiments considered faults located at three different positions relative to the load zone. The healthy and faulty bearings were reinstalled at both the drive end (DE) and fan end (FE) locations (where each configuration comprises either two healthy bearings or one healthy bearing and one faulty bearing), and data were collected synchronously, with one accelerometer at each location. For each configuration, experiments were made using four operational motor load conditions ranging from 0 (no load) to 3 horsepower (HP). In most cases, the experiments used a 12 kHz sampling rate, while some used 48 kHz. All the experiments consist of signals that are approximately 10 seconds long.
@@ 4.2.4 Summary
Dataset Acquisition locations Conditions per instance Acquistions per condition per location Total number of signals
UORED-VAFCLS 1 1 1 60
CWRU 2 4 1 144
Paderborn 1 4 20 1360
@@ 4.3.1 Shallow Learning models with bearing fault based features
In the work of [9], statistical metrics extracted from time-domain signals were employed, including mean, absolute median, standard deviation, skewness, kurtosis, crest factor, energy, RMS value, peak count, zero-crossing count, Shapiro-Wilk test statistic, and Kullback–Leibler divergence. In the frequency domain, [45] explored the use of bearing-specific fault frequencies, such as BPFI (Ball Pass Frequency of the Inner race), BPFO (Ball Pass Frequency of the Outer race), and BSF (Ball Spin Frequency). Both works reported favorable results using these handcrafted features, motivating their inclusion in our shallow learning pipeline. To highlight modulated fault components distributed across the spectrum, envelope analysis is performed to shift these components into the baseband, following the principles explained in [44]. Subsequently, algorithms are applied to extract the magnitudes of spectral peaks. In this study, the envelope spectrum is computed in the frequency range of 500 Hz to 10 kHz for the PU and UORED-VAFCLS datasets, which offer high sampling rates. For the CWRU dataset, limited to a 12 kHz sampling rate, a reduced band of 500 Hz to 6 kHz is adopted. Fault frequencies are computed based on the rotation speeds and geometric parameters provided within each dataset. Table 5 shows all features used on shallow learning models.
@@ 4.3.2 Deep Learning architectures
Informed by a review of contemporary deep learning models for vibration analysis, we selected the Wide First-layer Kernel Deep Convolutional Neural Network (WDCNN) [46] as the primary architecture for our investigation. This model, with approximately 172k parameters for an input segment length of 12,000 samples, offers a well-established baseline. To benchmark its performance against more recent or complex 1D convolutional architectures, a comparative analysis was conducted on the UORED-VAFCLS dataset, evaluating the WDCNN against the CDCN [47], WDTCNN [48], 1D-ConvNet [49], and RESNET1D [50] models. The results of this comparative study are presented in Section 5.1. To implement our multi-label classification framework, the output layer of each architecture was configured with a number of nodes equal to the number of detectable fault modes (excluding the healthy state). Each output node employs a sigmoid activation function, effectively operating as an independent binary classifier for a specific fault type.
@@
A critical challenge posed by the CWRU and UORED-VAFCLS datasets is the limited number of acquisitions, which substantially increases the risk of model overfitting. To mitigate this, we implemented a data augmentation strategy combining two techniques. The first, Random Crop, generates training samples by extracting segments of a predefined length (e.g., 42,000 samples, corresponding to 1 second for UORED-VAFCLS) from random positions within the full signal. The second, Random Gain (referred to in [51] as Scaling), introduces stochastic amplitude variations by multiplying each signal by a scalar drawn from a normal distribution $\mathcal{N}(\mu,\sigma^{2})$. To ensure sufficient gradient updates per epoch, we created virtual datasets by applying these augmentations, effectively increasing the number of training examples per epoch by a factor of 20 for UORED-VAFCLS and 5 for CWRU. These techniques are applied exclusively to the training set, whereas the test set is preserved as non-overlapping segmented signals, as described in the previous section.
@@
The augmentation hyperparameters were systematically optimized on the UORED-VAFCLS dataset. Following [51], the mean $\mu$ for the Random Gain distribution was fixed at 1.0, while the standard deviation $\sigma$ was tuned over the set 0.3, 0.5, 0.7. Our validation experiments revealed that the optimal configuration was a combination of Random Crop and Random Gain (RC+RG) with $\sigma=0.7$ for time-domain signals and $\sigma=0.3$ for frequency-domain inputs. To maintain a feasible computational budget, these optimized augmentation hyperparameters were subsequently applied without modification to all other datasets and experiments.
For the model selection phase (the inner loop of our CVM-CV protocol), we defined a hyperparameter search space to tune batch size, learning rate, and the input normalization strategy. Training was conducted for a fixed duration of 30 epochs across all experiments. The specific hyperparameter values explored during this tuning process are detailed in Table 6.

### v3

# arXiv:2509.22267v3[cs.LG] 03 Mar 2026

@@ 1 Introduction
Establishing that conventional (non-bearing-wise) splitting strategies indeed lead to unrealistic results—and thus should be avoided in future work—requires one to show that using a bearing-wise split causes a significant performance gap in an otherwise identical setup. This path was taken by [6, 7, 8, 9], which have observed accuracy dropping from near 100% to around 40%-60% depending on the specific setup. However, [7, 8] proposed splitting data by fault size, which does not necessarily corresponds to a bearing-wise split (see Section 3.4), while [7, 9] could not entirely eliminate leakage with respect to the healthy class in the CWRU dataset, as this dataset contains a single healthy bearing configuration.
@@
An alternative approach to completely eliminating data leakage is conducting inter-testbench experiments, typically framed within a cross-domain or domain generalization context. In this formulation, the testbenches used for training and testing are completely isolated; however, this introduces the significant challenge of domain generalization. We suspect that the bearing fault diagnosis literature has largely treated intra-domain classification as a solved problem, prompting a shift toward cross-domain research. Research by [10] indicates that many studies define a domain as a combination of operating conditions (e.g., load, rotation speed, torque) rather than distinct physical bearings. This perspective has led to various data splitting methodologies within single datasets, where data is partitioned according to these conditions. A limited number of studies have addressed the inter-testbench scenario by isolating datasets, such as [11], and [12]. Although these approaches effectively eliminate data leakage, domain generalization remains a complex challenge, as target datasets often exhibit disparate feature distributions, necessitating specialized techniques to extract domain-invariant features. The present work, however, focuses on a simpler yet fundamentally critical problem: training and testing within a single testbench.
@@
This paper aims to further advance the reliable development and deployment of ML models for bearing fault diagnosis. We propose a novel methodology that rigorously addresses data leakage and class imbalance in an intra-dataset scenario, particularly for datasets with limited healthy bearing data. Our approach formulates the problem as a binary multi-label classification for each sensor location (e.g., drive end and fan end), enabling the detection of the presence or absence of each fault type (e.g., inner, outer, ball). This formulation specifically addresses the disadvantages of multiclass accuracy, which serves as a poor proxy for real-world performance. Because accuracy treats all misclassifications as equal, it fails to distinguish between “false alarm” (False Positive) and a “missed detection” (False Negative), the latter being significantly more costly in industrial maintenance. Furthermore, in the presence of class imbalance, a model can achieve high accuracy by simply predicting the majority class, effectively “hiding” its inability to identify rare but critical fault states. Another issue caused by the multiclass formulation is the inability of detecting co-occurring faults. While it is possible to create classes for combined faults, it is often unpractical and most public datasets contain few examples of those cases. By treating faults as independent binary problems, we can accommodate co-occurring defects and utilize faulty signals from one label as true negatives for others—a significant advantage when healthy data is scarce, such as in the CWRU dataset. In this formulation, we can also utilize prevalence-independent metrics such as the Area Under the Receiver Operating Characteristic curve (AUROC). This allows for a more realistic representation of real-world conditions where decision thresholds must be precisely tuned to prioritize sensitivity or specificity based on specific safety and operational requirements.
@@
With these contributions, this work aims to foster the development of more robust and trustworthy ML models for bearing fault diagnosis, ensuring their performance more closely reflects their capabilities in real-world industrial settings. The source code for this paper can be found at github.com/gama-ufsc/bearing-data-leakage.
@@ 3.2 A Toy Example
To demonstrate the performance inflation caused by a non-bearing-wise data split, we constructed a synthetic binary fault detection experiment, in which the task is to classify each sample as either faulty or healthy. We simulate a feature space for a dataset of $B=48$ distinct bearings, which could be obtained, for example, as the output of a deep learning model’s feature extractor. This simulated space contains two feature types: 3 fault-predictive features and 48 bearing-identity features, resulting in 51 features. Each identity feature is unique to a specific bearing, representing a spurious correlation unrelated to the fault condition. All features were generated by adding zero-mean Gaussian noise with unit variance to a constant base value: $a_{f}=1.5$ for fault-predictive features and $a_{b}=8$ for bearing-identity features. All experiments were designed to maintain a fixed number of samples per bearing, using a base value of 40 samples each. The dataset comprised 24 healthy bearings and 24 faulty bearings, ensuring a consistently balanced class distribution.
@@ 3.3 Prevalence of data leakage in bearing fault diagnosis
As mentioned in the introduction, data leakage is widespread in the literature on bearing fault diagnosis, with previous works [4, 6] reporting data splitting issues in over 90% of published papers. To complement and update these findings, we conducted an investigation of papers published in 2025 using a similar methodology. First, we searched the Mechanical Systems and Signal Processing journal database using the queries “bearing fault diagnosis” and “machine learning”, which yielded 195 published papers from January 2025 up to December 2025. We then randomly selected 10 papers that focus on the problem of training and testing within a single dataset under a supervised machine learning context. Second, we extended our search to include papers from other journals within the same scope, compiling an additional 8 papers. Finally, we carefully analyzed the experimental methodology in each of these papers to understand how the data split was performed.
Table 1 lists these 18 papers, detailing whether they specified the train-test partitioning and, if so, how it was performed. We found that 8 of the 18 papers used a random splitting strategy, while 9 detailed a condition-wise split. The most common condition considered for partitioning was the load, although other papers considered rotation speed or level of noise. The remaining paper did not mention any partitioning methodology, which suggests that the authors may not have devoted sufficient attention to a detail critical for preventing leakage. Our conclusion is that despite an increase in the number of articles that detail their train-test partitioning methodology, data leakage remains a prevalent issue.
@@ 3.4 Related works
Hendriks et al. [7] proposed a fault-size splitting strategy for the CWRU dataset, considering it a better approach to obtain domain shift in bearing fault diagnosis. Since in the CWRU dataset each combination of fault size and type (i.e. inner, outer, ball) corresponds to a unique bearing, splitting by fault size indirectly produces as a bearing-wise partition. However, signals from the single healthy configuration (consisting of a healthy bearing in the fan end and another healthy bearing in the drive end) were included in both training and test sets, resulting in data leakage.
@@
A comprehensive investigation into dataset biases and evaluation protocols was conducted by [5], who proposed evaluation methodologies for widely used datasets such as CWRU, PU, MFPT, IMS, and UOC. Their approach focused on mitigating segmentation-level leakage, advocating for condition-wise splits across all datasets under a multiclass framework, with F1-macro reported as an auxiliary metric. However, the results remained overoptimistic due to residual bearing-level leakage, which was not addressed in their work.
Another notable contribution is the study by [6], which proposed a bearing-wise split under a multiclass classification setting, referred to as a “part-to-part” approach. Using the KAt and CMTH datasets, the authors demonstrated a high correlation between signals originating from the same bearing—violating the i.i.d. assumption—and showed the substantial drop in accuracy when shifting from a condition-wise to a bearing-wise split, revealing the effects of a bearing-level leakage. Although this paper propose a similar analysis to ours, their experiments contain confounding factors that make it difficult to understand the origin of the decrease in model performance. In their work, data splitting strategies such as “run-to-run”, “day-to-day” and “part-to-part” were proposed. Each strategy corresponds to completely different training set compositions, making it difficult to identify which one is more diverse. One could argue that the decrease in performance when changing to the “part-to-part” split is associated with the model being less diverse, rather than uniquely due to data leakage. Our paper focuses on solving this issue by proposing controlled data leakage experiments, where the training set remains fixed while varying the test sets.
@@ 4.2.2 Paderborn University (PU) Dataset
Crucially, prior work by [40] demonstrated that a significant domain shift exists between artificially damaged and naturally degraded bearings, leading to poor generalization when training on the former and testing on the latter. To circumvent this issue and ensure the practical relevance of our findings, our investigation exclusively utilizes the subset of bearings with naturally occurring faults from accelerated lifetime tests, along with the healthy reference bearings. Our analysis incorporates components K006 and KI17, thereby expanding upon the set used in the original benchmark study, which are shown in Table 3. Lastly, to limit computational overhead, all measurements in this dataset were resampled to 42 kHz. Although bearings with combined faults are in principle compatible with our multi-label methodology, their limited representation in the dataset hinders robust learning and evaluation. In fact, only three such bearings are available (IDs KB23, KB24, and KB27). Therefore, we exclude these samples from both training and testing to prevent biased or unreliable conclusions.
@@ 4.2.3 CWRU bearing fault dataset
The Case Western Reserve University (CWRU) bearing fault dataset is a collection of experiments that involved a single pair of healthy bearings and several artificially created faulty bearings [43]. The faults were created through electro-discharge machining, introducing point faults with diameters of 7, 14, 21, and 28 mils in the inner race, outer race, and rolling element separately. For the outer race faults, the experiments considered faults located at three different positions relative to the load zone. The healthy and faulty bearings were reinstalled at both the drive end (DE) and fan end (FE) locations (where each configuration comprises either two healthy bearings or one healthy bearing and one faulty bearing), and data were collected synchronously, with one accelerometer at each location. For each configuration, experiments were made using four operational motor load conditions ranging from 0 (no load) to 3 horsepower (HP). In most cases, the experiments used a 12 kHz sampling rate, while some used 48 kHz. All the experiments consist of signals that are approximately 10 seconds long.
@@ 4.2.4 Summary
Dataset Acquisition locations Conditions per instance acquisitions per condition per location Total number of signals Total number of signals in training
UORED-VAFCLS 1 1 1 60 36
CWRU 2 4 1 144 72
Paderborn 1 4 20 1360 880
@@ 4.3.1 Shallow Learning models with bearing fault based features
In the work of [9], statistical metrics extracted from time-domain signals were employed, including mean, absolute median, standard deviation, skewness, kurtosis, crest factor, energy, RMS value, peak count, zero-crossing count, Shapiro-Wilk test statistic, and Kullback–Leibler divergence. In the frequency domain, [44] explored the use of bearing-specific fault frequencies, such as BPFI (Ball Pass Frequency of the Inner race), BPFO (Ball Pass Frequency of the Outer race), and BSF (Ball Spin Frequency). Both works reported favorable results using these handcrafted features, which motivated the inclusion of a subset of them in our shallow learning pipeline. To highlight modulated fault components distributed across the spectrum, envelope analysis is performed to shift these components into the baseband, following the principles explained in [43]. Subsequently, algorithms are applied to extract the magnitudes of spectral peaks. In this study, the envelope spectrum is computed in the frequency range of 500 Hz to 10 kHz for the PU and UORED-VAFCLS datasets, which offer high sampling rates [^7] In [43], envelope analysis is performed without pre-filtering (wide-band analysis). Although the author notes that this approach is sufficient for the CWRU dataset, we restrict the bandwidth to exclude low frequencies (below 500 Hz) and very high frequencies (above 10 kHz), while still preserving a broad range that is potentially applicable to most signals. An inspection of the artificially damaged bearings on PU revealed that the majority of the relevant spectral content is concentrated within this interval.. For the CWRU dataset, limited to a 12 kHz sampling rate, a reduced band of 500 Hz to 6 kHz is adopted. Fault frequencies are computed based on the rotation speeds and geometric parameters provided within each dataset. Table 5 shows all features used on shallow learning models.
@@ 4.3.2 Deep Learning architectures
Informed by a review of contemporary deep learning models for vibration analysis, we selected the Wide First-layer Kernel Deep Convolutional Neural Network (WDCNN) [45] as the primary architecture for our investigation. This model, with approximately 172k parameters for an input segment length of 12,000 samples, offers a well-established baseline. To benchmark its performance against more recent or complex 1D convolutional architectures, a comparative analysis was conducted on the UORED-VAFCLS dataset, evaluating the WDCNN against the CDCN [46], WDTCNN [47], 1D-ConvNet [48], and RESNET1D [49] models. The results of this comparative study are presented in Section 5.1. To implement our multi-label classification framework, the output layer of each architecture was configured with a number of nodes equal to the number of detectable fault modes (excluding the healthy state). Each output node employs a sigmoid activation function, effectively operating as an independent binary classifier for a specific fault type. Model training was carried out using the Binary Cross-Entropy loss function, which is well suited for this multi-label setting.
@@
A critical challenge posed by the CWRU and UORED-VAFCLS datasets is the limited number of acquisitions, which substantially increases the risk of model overfitting. To mitigate this, we implemented a data augmentation strategy combining two techniques. The first, Random Crop, generates training samples by extracting segments of a predefined length (e.g., 42,000 samples, corresponding to 1 second for UORED-VAFCLS) from random positions within the full signal. The second, Random Gain (referred to in [50] as Scaling), introduces stochastic amplitude variations by multiplying each signal by a scalar drawn from a normal distribution $\mathcal{N}(\mu,\sigma^{2})$. These techniques are applied exclusively to the training set, whereas the test set is preserved as non-overlapping segmented signals, as described in the previous section.
@@
The augmentation hyperparameters were systematically optimized on the UORED-VAFCLS dataset using the WDCNN architecture. Following [50], the mean $\mu$ for the Random Gain distribution was fixed at 1.0, while the standard deviation $\sigma$ was tuned over the set 0.3, 0.5, 0.7. Our validation experiments revealed that the optimal configuration was a combination of Random Crop and Random Gain (RC+RG) with $\sigma=0.7$ for time-domain signals and $\sigma=0.3$ for frequency-domain and spectrum envelope inputs. To maintain a feasible computational budget, these optimized augmentation hyperparameters were subsequently applied without modification to all other datasets and experiments.
For the model selection phase (the inner loop of our CVM-CV protocol), we defined a hyperparameter search space to tune batch size, learning rate, and the input normalization strategy, which is detailed in Table 6. We considered standard scaling normalization under two approaches: entry-wise and global. Entry-wise normalization scales each signal using its own mean and standard deviation, whereas global normalization uses the mean and standard deviation computed over the training set. Training was conducted for a fixed duration of 600 epochs across all experiments on the UORED-VAFCLS dataset, which contains 36 training signals under our splitting strategy. This number of epochs was chosen to ensure sufficient gradient updates to obtain stable validation performance curves, while the data augmentation strategies were used to help the curves to reach a plateau without subsequently decreasing, avoiding the need for early stopping. Due to their larger sizes, the CWRU (72 training signals) and PU (880 training signals) datasets required 150 and 30 epochs, respectively, to reach a similar plateau on the validation curves.